OpenStudy (anonymous):

2(5x+20)=2(6x+12) What is x?

OpenStudy (anonymous):

5x +20 = 6x +12 20 = x +12 8 = x

OpenStudy (anonymous):

Thanks

OpenStudy (saifoo.khan):

Cancel the 2s on the both side U will be left with: 5x + 20 = 6x + 12 Shuffle: x = 8

OpenStudy (anonymous):

Gotcha!

OpenStudy (anonymous):

x = 7/4

OpenStudy (saifoo.khan):

Hi there.

OpenStudy (anonymous):

hi boo

OpenStudy (saifoo.khan):

how's your son? and how was your final?